# D-inverse constellations

### v1

@@ 3.2. D-inverses
In a constellation $\mathcal{Q}$, every $e\in D(\mathcal{Q})$ is a D-inverse for itself, and each element of $Q$ has at most one D-inverse if and only if $\mathcal{Q}$ is normal.
@@
For $e\in D(\mathcal{Q})$, $e\cdot e=e=D(e)$ so $e$ is a D-inverse of itself.
Suppose $\mathcal{Q}$ is normal and $s_{1}$ and $s_{2}$ are both D-inverses of $s\in Q$. So $s_{1}\cdot s$ and $s\cdot s_{1}$ exist, whence so does $s_{1}\cdot(s\cdot s_{1})=s_{1}\cdot D(s)=s_{1}$ by (Const2). So by (Const1),
@@
Conversely, if D-inverses are unique, suppose $e\cdot f$ and $f\cdot e$ both exist where $e,f\in D(\mathcal{Q})$. Then $e\cdot f=D(e)$ and $f\cdot e=D(f)$, and so $f$ is a D-inverse of $e$, whence $f=e$ by uniqueness. This establishes normality. ∎
@@
We say the constellation $\mathcal{Q}$ is a D-inverse constellation if every element of $Q$ has a unique D-inverse. We use $s^{\prime}$ to denote the unique D-inverse of $s$.
@@
The constellation $\mathcal{Q}$ is D-inverse if and only if every element of $Q$ has a D-inverse and $\mathcal{Q}$ is normal.
@@
Suppose $\mathcal{P},\mathcal{Q}$ are constellations, with $\mathcal{P}$ being D-inverse and $\mathcal{Q}$ normal, and suppose that $\rho:\mathcal{P}\rightarrow\mathcal{Q}$ is a surjective radiant. Then $\mathcal{Q}$ is D-inverse, and $(s\rho)^{\prime}=s^{\prime}\rho$.
@@
A monoid is a D-inverse constellation if and only if it is a group. Every poset $P$ gives a D-inverse constellation if we define $e\cdot f=e$ if and only if $e\leq f$; this is a constellation in which $D(e)=e$ for all $e\in P$, as noted first in [1], and normal as noted in [3], and since $e\cdot e=e=D(e)$ for all $e\in P$, we have that $e$ is a D-inverse of itself.
A further significant example of a D-inverse constellation is the constellation $\mathcal{I}_{X}$ of one-to-one partial functions on the set $X$. The constellation $\mathcal{I}_{X}$ comes equipped with the additional operation of inversion: this is because for all $f\in I_{X}$, if $f^{\prime}$ is the inverse of $f$ in the usual partial function sense, then $f\cdot f^{\prime}=D(f)$ and $f^{\prime}\cdot f=D(f^{\prime})$, so $f^{\prime}$ is the D-inverse of $f$ in the constellation sense also. The inverse semigroup $\mathscr{I}_{X}$ and the D-inverse constellation $\mathcal{I}_{X}$ have the same underlying set, but $f\cdot g$ exists in $\mathscr{I}_{X}$ if and only if $fD(g)=fgg^{\prime}=f$ in $\mathscr{I}_{X}$. In fact this generalises as follows, courtesy of the Vagner-Preston representation theorem, which says that any inverse semigroup embeds as a subsemigroup of some $\mathscr{I}_{X}$, (although a direct proof is straightforward).
@@
so that $\rho_{s}\cdot\rho_{s^{\prime}}=\rho_{D(s)}=D(s\rho)$. Similarly, $\rho_{s^{\prime}}\cdot\rho_{s}=D({s^{\prime}}\rho)$. All of this shows that $s^{\prime}\rho=\rho_{s^{\prime}}=(\rho_{s})^{\prime}=(s\rho)^{\prime}$, as required. ∎
From (Const2), in a D-inverse constellation $\mathcal{Q}$ we have $s\cdot(s^{\prime}\cdot s)=(s\cdot s^{\prime})\cdot s=s$ for all $s\in Q$. Other familiar facts of inverse semigroup theory such as $(st)^{\prime}=t^{\prime}s^{\prime}$ do not carry over since the existence of $s\cdot t$ does not ensure that $t^{\prime}\cdot s^{\prime}$ exists (even if both $s,t$ have D-inverses). For example, let $X=\{x,y\}$, with $s,t\in I_{X}$ defined as follows: $s=\{(x,x)\}$ and $t=1$ (the identity function on $X$). Then in $\mathcal{I}_{X}$, $s^{\prime}=s$, $t^{\prime}=t$, $s\cdot t=s$ yet $t^{\prime}\cdot s^{\prime}=t\cdot s$ does not exist. However, we do have the following.
@@ 4. D-inverse constellations are exactly ordered groupoids
In the examples of constellations considered in [3], most had a notion of range, satisfying some natural properties. In [4], it was shown that they are nothing but ordered categories with restriction, a class containing all ordered groupoids. In this section we show that every D-inverse constellation is a constellation with range, and that when viewed as ordered categories with restriction, they are nothing but ordered groupoids.
@@ 4.1. Constellations with range and ordered categories with restrictions
$s_{D}=\{e\in D(\mathcal{Q})\mid s\cdot e\mbox{ exists, and for all }f\in D(\mathcal{Q}),s\cdot f\mbox{ exists implies }e\leq f\}.$
@@
[3] Suppose $\mathcal{Q}=(Q,\,\cdot\,,D,R)$ is a constellation with pre-range. Let $s,t\in Q$. Then the following are equivalent: $s\cdot t$ exists; $s\cdot D(t)$ exists; $R(s)\cdot D(t)$ exists; $R(s)\cdot t$ exists. Further, if $s\cdot t$ exists, then $R(s\cdot t)\leq R(t)$.
@@
The above definition coincides with the usual definition of “epimorphism” if $\mathcal{Q}$ arises from a category. It is also reminiscent of the condition that an element $a$ of a semigroup $S$ be $\mathcal{L}^{*}$-related to an idempotent $R(a)$. Here $\mathcal{L}^{*}$ is the relation of mutual cancellativity given by $a\,\mathcal{L}^{*}\,b$ if for any $x,y\in S^{1}$, we have $ax=ay$ if and only if $bx=by$. The constellation with range $\mathcal{C}_{X}$ is left cancellative.
@@
If $\mathcal{Q}=(Q,\,\cdot\,,D,R)$ is a (left cancellative) constellation with range, then the derived category $(Q,\,\circ\,,D,R)$, where $s\circ t:=s\cdot t$ is defined if and only if $R(s)=D(t)$, is an ordered (left cancellative) category with restrictions in which $\leq$ is the natural order on $\mathcal{Q}$ as a normal constellation, and for $e\leq D(s)$, $e|s=e\cdot s$.
Conversely, if $\mathcal{C}=(C,\,\circ\,,D,R,\leq)$ is a (left cancellative) ordered category with restrictions, then setting $s\cdot t$ equal to $s\circ(R(s)|t)$ whenever $R(s)\leq D(t)$ makes $(C,\,\cdot\,,D,R)$ into a (left cancellative) constellation with range, and the given partial order is nothing but the natural order on the constellation.
With the morphisms between ordered categories with restriction taken to be order-preserving functors, we obtain a category ${\mathcal{C}}$. We denote by ${\mathcal{Q}}$ the category of constellations with range and range radiants. The natural extension of $\mathbf{C}$ and $\mathbf{Q}$ to morphisms yield functors between these categories.
@@
[4, Theorem 5.10] The categories ${\mathcal{C}}$ and ${\mathcal{Q}}$ are isomorphic. Moreover, ${\bf C}({\bf Q}({\mathcal{C}}))={\mathcal{C}}$ and ${\bf Q}({\bf C}({\mathcal{Q}}))={\mathcal{Q}}$.
@@ 4.2. The D-inverse constellation case
Every D-inverse constellation is a constellation with range in which $R(s)=s^{\prime}\cdot s=D(s^{\prime})$ for all $s\in\mathcal{Q}$, is right cancellative and is left cancellative as a constellation with range.
@@
Let $\mathcal{Q}$ be a D-inverse constellation and define $R(s)$ as in the statement of the proposition, for any $s\in\mathcal{Q}$. Now $R(s)=s^{\prime}\cdot s=D(s^{\prime})\in D(\mathcal{Q})$ and $s\cdot R(s)=s\cdot(s^{\prime}\cdot s)=(s\cdot s^{\prime})\cdot s=D(s)\cdot s=s$. Suppose $s\cdot e$ exists for some $e\in D(S)$. Then $s^{\prime}\cdot(s\cdot e)=(s^{\prime}\cdot s)\cdot e=R(s)\cdot e$ exists, and so $R(s)\leq e$. Since $\mathcal{Q}$ is normal, $\leq$ is a partial order, and so $\mathcal{Q}$ is a constellation with pre-range.
If $s\cdot t$ (equivalently by Result 4.2, $R(s)\cdot t$) exists, then $s^{\prime}\cdot(s\cdot t)=(s^{\prime}\cdot s)\cdot t$ exists. Hence by Result 4.2,
@@
$=R((s^{\prime}\cdot s)\cdot t)=R(s^{\prime}\cdot(s\cdot t))\leq R(s\cdot t).$
@@
Suppose $s\in\mathcal{Q}$ and $e,f\in D(\mathcal{Q})$ are such that $R(e\cdot s)=R(f\cdot s)$. Then $(e\cdot s)^{\prime}\cdot(e\cdot s)=(f\cdot s)^{\prime}\cdot(f\cdot s)$. But $s\cdot s^{\prime}=D(s)$ exists, and so $(e\cdot s)\cdot s^{\prime}=e\cdot(s\cdot s^{\prime})=e\cdot D(s)=e$; similarly, $(f\cdot s)\cdot s^{\prime}=f$. Hence $(e\cdot s)^{\prime}\cdot((e\cdot s)\cdot s^{\prime})=(e\cdot s)^{\prime}\cdot e=(e\cdot s)^{\prime}$ and similarly $(f\cdot s)^{\prime}\cdot((f\cdot s)\cdot s^{\prime})=(f\cdot s)^{\prime}$. So
@@
It follows that $e\cdot s=f\cdot s$, and by the right cancellative property, $e=f$. So $\mathcal{Q}$ is strongly right cancellative as a constellation with range. ∎
@@
Let $X=\{x,y,z\}$ and $P=\{f,i,g,a,b\}\subseteq\mathcal{C}_{X}$, with $f=\{(x,x)\}$, $i=\{(x,x),(y,y)\}$, $g=\{(z,z)\}$, $a=\{(x,z)\}$ and $b=\{(x,z),(y,z)\}$. It is easy to see that $(P,\cdot,D,R)$ is a subconstellation of $(\mathcal{C}_{X},\,\cdot\,,D,R)$ that is also closed under range and hence is a constellation with range itself, moreover one that satisfies the congruence condition since $\mathcal{C}_{X}$ does. Moreover if $s\cdot u=t\cdot u$ for any $s,t,u\in P$, if $u\in D(P)$ then obviously $s=t$, but if not then $u\in\{a,b\}$ and $s,t\in D(P)$ and so $s=D(s)=D(s\cdot u)=D(t\cdot u)=D(t)=t$. However, $R(f\cdot b)=R(a)=g=R(b)=R(i\cdot b)$, yet $f\neq i$. So even for constellations with range satisfying the congruence condition, the second part of the definition of being strongly right cancellative is independent of the right cancellative law.
@@
Let $\mathcal{Q}$ be a D-regular constellation with range. Then $\mathcal{Q}$ is a D-inverse constellation, and $R(s)=D(s^{\prime})$.
@@
So $s\cdot t^{\prime}=s\cdot(R(s)\cdot t)=(s\cdot R(s))\cdot t=s\cdot t=D(s)$, and $t^{\prime}\cdot s=R(s)$, so $D(t^{\prime})=D(t^{\prime}\cdot s)=R(s)$. Hence $t^{\prime}$ is an (hence the) D-inverse of $s$. Moreover, $\mathcal{Q}$ is normal by Result 4.3. ∎
We have that $\mathcal{C}_{X}$ is a constellation with range, but it is not D-regular (since it is not even right cancellative as a constellation, as easy examples show). However, it is not hard to check that for all $a\in\mathcal{C}_{X}$ there exists $b\in\mathcal{C}_{X}$ for which $b\cdot a=R(a)$.
@@
The categories of D-inverse constellations and ordered groupoids are isomorphic, via the restriction of the functors $\bf C$ and $\bf Q$.
@@
Suppose $\mathcal{Q}$ is a D-inverse constellation. Then viewing it as a constellation with range, ${\bf C}(\mathcal{Q})$ is an ordered category with restrictions. Now $D(s)=s\cdot s^{\prime}$ and $D(s^{\prime})=s^{\prime}\cdot s$ in $\mathcal{Q}$, where $s^{\prime}$ is the D-inverse of $s$ in $\mathcal{Q}$. But $R(s)=D(s^{\prime})$ and $R(s^{\prime})=D(s^{\prime\prime})=D(s)$ by uniqueness, so $s\cdot s^{\prime}=s\circ s^{\prime}$ and $s^{\prime}\cdot s=s^{\prime}\circ s$ in ${\bf C}(\mathcal{Q})$.
It remains to show that (OG1) holds. Suppose $s,t\in{\bf C}(\mathcal{Q})$, with $s^{\prime},t^{\prime}$ the D-inverses of $s,t$ in $\mathcal{Q}$. Suppose that $s\leq t$. Then $s=D(s)|t$ with $D(s)\leq D(t)$, so $s=D(s)\cdot t=(s\cdot s^{\prime})\cdot t$ in $\mathcal{Q}$. As $D(s)\leq D(t)$, so $(s\cdot s^{\prime})\cdot(t\cdot t^{\prime})$ exists and equals $s\cdot s^{\prime}$. Then
@@
For the converse, suppose $\mathcal{C}$ is an ordered groupoid. Then it is an ordered category with restrictions, so that ${\bf Q}(\mathcal{C})$ is a constellation with range. Moreover for $s\in{\bf Q}(\mathcal{C})$, if $s^{\prime}$ is its inverse in $C$, then $D(s)=R(s^{\prime})$, so in $\mathcal{C}$ and hence in ${\bf Q}(\mathcal{C})$, $s\cdot s^{\prime}=s\circ s^{\prime}=D(s)$ and $s^{\prime}\cdot s=s^{\prime}\circ s=R(s)=D(s^{\prime})$, and so $s^{\prime}$ is a D-inverse of $s$ in ${\bf Q}(\mathcal{C})$, unique by normality (Proposition 3.9). Hence ${\bf Q}(\mathcal{C})$ is a D-inverse constellation.
The remainder of the details of the correspondence now follow from Result 4.7, also using the comment before Proposition 3.12. ∎

### v2

@@ 3.2. D-inverses
In a constellation $\mathcal{Q}$, every $e\in D(\mathcal{Q})$ is a D-inverse for itself, and each element of $\mathcal{Q}$ has at most one D-inverse if and only if $\mathcal{Q}$ is normal.
@@
[…]
Suppose that $\mathcal{Q}$ is normal and $s_{1}$ and $s_{2}$ are both D-inverses of $s\in Q$. Then $s_{1}\cdot s$ and $s\cdot s_{1}$ exist, whence so does $s_{1}\cdot(s\cdot s_{1})=s_{1}\cdot D(s)=s_{1}$ by (Const2). By (Const1),
@@
Conversely, if D-inverses are unique, suppose that $e\cdot f$ and $f\cdot e$ both exist where $e,f\in D(\mathcal{Q})$. Then $e\cdot f=e=D(e)$ and $f\cdot e=f=D(f)$, and so $f$ is a D-inverse of $e$, whence $f=e$ by uniqueness. This establishes normality. ∎
@@
We say the constellation $\mathcal{Q}$ is a D-inverse constellation if every element of $\mathcal{Q}$ has a unique D-inverse. We use $s^{\prime}$ to denote the unique D-inverse of $s$.
@@
The constellation $\mathcal{Q}$ is D-inverse if and only if every element of $\mathcal{Q}$ has a D-inverse and $\mathcal{Q}$ is normal.
@@
Suppose that $\mathcal{P},\mathcal{Q}$ are constellations, with $\mathcal{P}$ being D-inverse and $\mathcal{Q}$ normal, and suppose that $\rho:\mathcal{P}\rightarrow\mathcal{Q}$ is a surjective radiant. Then $\mathcal{Q}$ is D-inverse, and $(s\rho)^{\prime}=s^{\prime}\rho$.
@@
A monoid is a D-inverse constellation if and only if it is a group. Every poset $P$ gives a D-inverse constellation if we define $e\cdot f=e$ if and only if $e\leq f$; this is a constellation in which $D(e)=e$ for all $e\in P$, as noted first in [1], and normal as noted in [3]. Since $e\cdot e=e=D(e)$ for all $e\in P$, we have that $e$ is a D-inverse of itself.
A further significant example of a D-inverse constellation is the constellation $\mathcal{I}_{X}$ of one-to-one partial functions on the set $X$. The constellation $\mathcal{I}_{X}$ comes equipped with the additional operation of inversion: this is because for all $f\in\mathcal{I}_{X}$, if $f^{\prime}$ is the inverse of $f$ in the usual partial function sense, then $f\cdot f^{\prime}=D(f)$ and $f^{\prime}\cdot f=D(f^{\prime})$, so $f^{\prime}$ is the D-inverse of $f$ in the constellation sense also. The inverse semigroup $\mathscr{I}_{X}$ and the D-inverse constellation $\mathcal{I}_{X}$ have the same underlying set, but $f\cdot g$ exists in $\mathcal{I}_{X}$ if and only if $fD(g)=fgg^{\prime}=f$ in $\mathscr{I}_{X}$. In fact this generalises as follows, courtesy of the Vagner-Preston representation theorem, which says that any inverse semigroup embeds as a subsemigroup of some $\mathscr{I}_{X}$, (although a direct proof is straightforward).
@@
so that $\rho_{s}\cdot\rho_{s^{\prime}}=\rho_{D(s)}=D(s\rho)$. Similarly, $\rho_{s^{\prime}}\cdot\rho_{s}=D({s^{\prime}}\rho)$. Together, these statements show that $s^{\prime}\rho=\rho_{s^{\prime}}=(\rho_{s})^{\prime}=(s\rho)^{\prime}$, as required. ∎
From (Const2), in a D-inverse constellation $\mathcal{Q}$ we have $s\cdot(s^{\prime}\cdot s)=(s\cdot s^{\prime})\cdot s=s$ for all $s\in Q$. Other familiar facts of inverse semigroup theory such as $(st)^{\prime}=t^{\prime}s^{\prime}$ do not carry over since the existence of $s\cdot t$ does not ensure that $t^{\prime}\cdot s^{\prime}$ exists (even if both $s,t$ have D-inverses). For example, let $X=\{x,y\}$, with $s,t\in\mathcal{I}_{X}$ defined as follows: $s=\{(x,x)\}$ and $t=1$ (the identity function on $X$). Then in $\mathcal{I}_{X}$, $s^{\prime}=s$, $t^{\prime}=t$, $s\cdot t=s$ yet $t^{\prime}\cdot s^{\prime}=t\cdot s$ does not exist. However, we do have the following.
@@ 4. D-inverse constellations are exactly ordered groupoids
In the examples of constellations considered in [3], most had a notion of range, satisfying some natural properties. In [4], it was shown that they are nothing but ordered categories with restriction, a class containing all ordered groupoids. In this section we show that every D-inverse constellation is a constellation with range, and that when viewed as an ordered category with restriction, it is nothing but an ordered groupoid.
@@ 4.1. Constellations with range and ordered categories with restrictions
[…]
@@
[3] Suppose that $\mathcal{Q}=(Q,\,\cdot\,,D,R)$ is a constellation with pre-range. Let $s,t\in Q$. Then the following are equivalent: $s\cdot t$ exists; $s\cdot D(t)$ exists; $R(s)\cdot D(t)$ exists; $R(s)\cdot t$ exists. Further, if $s\cdot t$ exists, then $R(s\cdot t)\leq R(t)$.
@@
The above definition coincides with the usual definition of “epimorphism” if $\mathcal{Q}$ arises from a category. It is also reminiscent of the condition that an element $a$ of a semigroup $S$ be $\mathcal{L}^{*}$-related to an idempotent $R(a)$. Here $\mathcal{L}^{*}$ is the relation given by $a\,\mathcal{L}^{*}\,b$ if for any $x,y\in S^{1}$, we have $ax=ay$ if and only if $bx=by$. The constellation with range $\mathcal{C}_{X}$ is left cancellative.
@@
If $\mathcal{Q}=(Q,\,\cdot\,,D,R)$ is a ((left cancellative)) constellation with range, then the derived category $(Q,\,\circ\,,D,R)$, where $s\circ t:=s\cdot t$ is defined if and only if $R(s)=D(t)$, is an ordered ((left cancellative)) category with restrictions in which $\leq$ is the natural order on $\mathcal{Q}$ as a normal constellation, and for $e\leq D(s)$, we have that $e|s=e\cdot s$.
Conversely, if $\mathcal{C}=(C,\,\circ\,,D,R,\leq)$ is a ((left cancellative)) ordered category with restrictions, then setting $s\cdot t$ equal to $s\circ(R(s)|t)$ whenever $R(s)\leq D(t)$ makes $(C,\,\cdot\,,D,R)$ into a ((left cancellative)) constellation with range, and the given partial order is nothing but the natural order on the constellation.
If $\mathcal{Q}=(Q,\cdot\,,D,R)$ is a constellation with range and $\mathcal{C}=(C,\circ,D,R,\leq)$ is an ordered category with restrictions, denote by ${\bf C}(\mathcal{Q})$ the ordered category with restrictions $(Q,\circ,D,R,\leq)$ obtained as in Result 4.6 from $\mathcal{Q}$, and denote by ${\bf Q}(\mathcal{C})$ the constellation with range $(C,\circ,D,R)$ obtained as in that result from the ordered category with restrictions $\mathcal{C}$}. These two constructions are mutually inverse.
@@
Let $\mathcal{C}$} be an ordered category with restrictions and $\mathcal{Q}$} a constellation with range. Then ${\bf C}({\bf Q}({\mathcal{C}}))={\mathcal{C}}$ and ${\bf Q}({\bf C}({\mathcal{Q}}))={\mathcal{Q}}$.
The class of constellations with range is a category if arrows are taken to be range radiants, and the class of ordered categories with restrictions is itself a category in which the arrows are functors $\rho$ that are order-preserving, which means that $s\leq t$ implies $s\rho\leq t\rho$.
Result 4.8.
The category with objects ordered categories with restrictions and arrows being order preserving radiants is isomorphic to the category with objects constellations with range and arrows being range radiants.
@@ 4.2. The D-inverse constellation case
Every D-inverse constellation $\mathcal{Q}$ is a constellation with range in which $R(s)=s^{\prime}\cdot s=D(s^{\prime})$ for all $s\in\mathcal{Q}$, which is right cancellative, and left cancellative as a constellation with range.
@@
Let $\mathcal{Q}$ be a D-inverse constellation and define $R(s)$ as in the statement of the proposition, for each $s\in\mathcal{Q}$. Now $R(s)=s^{\prime}\cdot s=D(s^{\prime})\in D(\mathcal{Q})$ and $s\cdot R(s)=s\cdot(s^{\prime}\cdot s)=(s\cdot s^{\prime})\cdot s=D(s)\cdot s=s$. Suppose that $s\cdot e$ exists for some $e\in D(S)$. Then $s^{\prime}\cdot(s\cdot e)=(s^{\prime}\cdot s)\cdot e=R(s)\cdot e$ exists, and so $R(s)\leq e$. Since $\mathcal{Q}$ is normal, $\leq$ is a partial order, and so $\mathcal{Q}$ is a constellation with pre-range.
If $s\cdot t$ (equivalently, by Result 4.2, $R(s)\cdot t$) exists, then $s^{\prime}\cdot(s\cdot t)=(s^{\prime}\cdot s)\cdot t$ exists. Hence by Result 4.2,
@@
$R((s^{\prime}\cdot s)\cdot t)=R(s^{\prime}\cdot(s\cdot t))\leq R(s\cdot t).$
@@
Suppose that $s\in\mathcal{Q}$ and $e,f\in D(\mathcal{Q})$ are such that $R(e\cdot s)=R(f\cdot s)$. Then $(e\cdot s)^{\prime}\cdot(e\cdot s)=(f\cdot s)^{\prime}\cdot(f\cdot s)$. But $s\cdot s^{\prime}=D(s)$ exists, and so $(e\cdot s)\cdot s^{\prime}=e\cdot(s\cdot s^{\prime})=e\cdot D(s)=e$; similarly, $(f\cdot s)\cdot s^{\prime}=f$. Hence $(e\cdot s)^{\prime}\cdot((e\cdot s)\cdot s^{\prime})=(e\cdot s)^{\prime}\cdot e=(e\cdot s)^{\prime}$ and similarly $(f\cdot s)^{\prime}\cdot((f\cdot s)\cdot s^{\prime})=(f\cdot s)^{\prime}$. So
@@
It follows that $e\cdot s=f\cdot s$ and, by the right cancellative property, $e=f$. So $\mathcal{Q}$ is strongly right cancellative as a constellation with range. ∎
@@
Let $X=\{x,y,z\}$ and $P=\{f,i,g,a,b\}\subseteq C_{X}$, with $f=\{(x,x)\}$, $i=\{(x,x),(y,y)\}$, $g=\{(z,z)\}$, $a=\{(x,z)\}$ and $b=\{(x,z),(y,z)\}$. It is easy to see that $\mathcal{P}=(P,\cdot,D,R)$ is a subconstellation of $\mathcal{C}_{X}=(C_{X},\,\cdot\,,D,R)$ that is also closed under range and hence is a constellation with range itself, moreover one that satisfies the congruence condition since $\mathcal{C}_{X}$ does. Moreover if $s\cdot u=t\cdot u$ for any $s,t,u\in P$, if $u\in D(P)$ then obviously $s=t$, but if not then $u\in\{a,b\}$ and $s,t\in D(P)$ and so $s=D(s)=D(s\cdot u)=D(t\cdot u)=D(t)=t$. However, $R(f\cdot b)=R(a)=g=R(b)=R(i\cdot b)$, yet $f\neq i$. So even for constellations with range satisfying the congruence condition, the second part of the definition of being strongly right cancellative is independent of the right cancellative law.
@@
Let $\mathcal{Q}$ be a D-regular constellation with range. Then $\mathcal{Q}$ is a D-inverse constellation, and $R(s)=D(s^{\prime})$ for all $s\in\mathcal{Q}$.
@@
We have that $s\cdot t^{\prime}=s\cdot(R(s)\cdot t)=(s\cdot R(s))\cdot t=s\cdot t=D(s)$ and, as just shown, $t^{\prime}\cdot s=R(s)$, so $D(t^{\prime})=D(t^{\prime}\cdot s)=R(s)$. Hence $t^{\prime}$ is a D-inverse of $s$, so $\mathcal{Q}$ is D-regular. Moreover, $\mathcal{Q}$ is normal by Result 4.3, and so by Corollary 3.10, $\mathcal{Q}$ is D-inverse. ∎
We have that $\mathcal{C}_{X}$ is a constellation with range, but it is not D-regular (since it is not even right cancellative as a constellation, as easy examples show). However, it is not hard to check that for all $a\in C_{X}$ there exists $b\in C_{X}$ for which $b\cdot a=R(a)$.
@@
The category with objects D-inverse constellations and radiants as arrows is isomorphic to the category with objects ordered groupoids and order-preserving functors as arrows, via the restriction of the functors as in Result 4.8.
@@
Suppose that $\mathcal{Q}$ is a D-inverse constellation. Then viewing it as a constellation with range, ${\bf C}(\mathcal{Q})$ is an ordered category with restrictions. Now $D(s)=s\cdot s^{\prime}$ and $D(s^{\prime})=s^{\prime}\cdot s$ in $\mathcal{Q}$, where $s^{\prime}$ is the D-inverse of $s$ in $\mathcal{Q}$. But $R(s)=D(s^{\prime})$ and $R(s^{\prime})=D(s^{\prime\prime})=D(s)$ by uniqueness, so $s\cdot s^{\prime}=s\circ s^{\prime}$ and $s^{\prime}\cdot s=s^{\prime}\circ s$ in ${\bf C}(\mathcal{Q})$.
It remains to show that (OG1) holds. Suppose that $s,t\in{\bf C}(\mathcal{Q})$, with $s^{\prime},t^{\prime}$ being the D-inverses of $s,t$ in $\mathcal{Q}$. Suppose that $s\leq t$. Then $s=D(s)|t$ with $D(s)\leq D(t)$, so $s=D(s)\cdot t=(s\cdot s^{\prime})\cdot t$ in $\mathcal{Q}$. As $D(s)\leq D(t)$, so $(s\cdot s^{\prime})\cdot(t\cdot t^{\prime})$ exists and equals $s\cdot s^{\prime}$. Then
@@
For the converse, suppose that $\mathcal{C}$ is an ordered groupoid. Then it is an ordered category with restrictions, so that ${\bf Q}(\mathcal{C})$ is a constellation with range. Moreover for $s\in{\bf Q}(\mathcal{C})$, if $s^{\prime}$ is its inverse in $C$, then $D(s)=R(s^{\prime})$, so in $\mathcal{C}$ and hence in ${\bf Q}(\mathcal{C})$, $s\cdot s^{\prime}=s\circ s^{\prime}=D(s)$ and $s^{\prime}\cdot s=s^{\prime}\circ s=R(s)=D(s^{\prime})$, and so $s^{\prime}$ is a D-inverse of $s$ in ${\bf Q}(\mathcal{C})$, which is unique by normality (Proposition 3.8). Hence ${\bf Q}(\mathcal{C})$ is a D-inverse constellation.
The remainder of the details of the correspondence now follow from Result 4.8, also using the comment before Proposition 3.11. ∎
@@
Starting with an inverse semigroup $S$, we may view it as a D-inverse constellation by setting $D(s)=ss^{\prime}$ and $s\cdot t=st$ but only when $stt^{\prime}=s$, as in Proposition 3.12, and then this may be viewed as an ordered groupoid as in Theorem 4.14 by retaining $D(s)$, setting $R(s)=D(s^{\prime})=s^{\prime}s$, setting $s\circ t=st$ whenever $R(s)=D(t)$, which is to say that $s^{\prime}s=tt^{\prime}$, and the order is given by the natural order, namely $s\leq t$ if and only if $s=D(s)t=ss^{\prime}t$. On the other hand, beginning with the same inverse semigroup $S$, the usual way to view it as an inductive (hence ordered) groupoid is to define $D(s)=ss^{\prime}$, $R(s)=s^{\prime}s$, $s\circ t=st$ when $s^{\prime}s=R(s)=D(t)=tt^{\prime}$, and the order given by $s\leq t$ if $s=ss^{\prime}t$. Clearly, these two viewpoints coincide.